News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Wednesday 20 October 1999 11.00am EST

During the last 24 hours Chandra completed observations of the GTO targets Cygnus X-1 and Cygnus X-3 with ACIS-S and HETG. During the next 24 hours observations will be conducted of PKS 2155-304 and 4U1636-53 with ACIS-S and the HETG, and an observation started of the target A697 with ACIS-I.

The re-built load containing re-configured ACIS parameter blocks to avoid use of FEP-0 was uplinked during the evening of 10/19/99 and is operating as expected. This load will continue through 12AM 10/24 when the next mission schedule will start. The products for this schedule are expected to be in review today.

The spacecraft systems continued to operate nominally.

Return to Status Reports